# Citation Sentiment Reflects Multiscale Sociocultural Norms

### v1

@@ V.3 Departmental and Cultural Measures
Alongside citation sentiment, we evaluate other measures at the broader departmental and cultural scales. At the departmental level, we evaluate the proportion of articles originating in benchwork, terming this measure the benchwork score. Lower values correspond to computational settings, whereas larger values correspond to research performed predominantly in wet labs. We measure this score using a base version (un-finetuned) of GPT-3.5-Turbo-1106. For each of the $28$ departments, we first sample $100$ articles, excluding reviews. We then prompt the model to determine whether a given article is a product of benchwork by examining its abstract. Abstracts are enclosed within <abstract> tags in the PMC OA XML data. The benchwork score for a department is the fraction of papers identified by the model as having originated in a wet lab. The prompt text supplied to the LLM is available in the Supplement. We also measure synthesis as the proportion of review papers among all papers in a discipline. We identify a review paper by the article-type variable from the XML file for each paper. At the broader cultural scale, we use the 6-D model of national culture to measure power distance and individualism. Power distance is the “extent to which the less powerful members of organizations and institutions (like the family) accept and expect that power is distributed unequally.” Individualism, in turn, is the “extent to which people feel independent, as opposed to being interdependent as members of larger wholes.” Values for these measures are available at [74, 61].
@@ Acknowledgements.
The authors gratefully acknowledge helpful discussions with Drs. Jordan D. Dworkin, Julia K. Brynildsen, and David Lydon-Staley.

### v2

@@ V.3 Departmental and Cultural Measures
Alongside citation sentiment, we evaluated other measures at the broader departmental and cultural scales. To examine the disciplinary impact on sentiment, we selected $28$ departments (see Supplement for more details). For the corresponding regression analyses of disciplinary and global scales, we included only the subsets that have at least $100$ citations towards collaborators to ensure sufficient sample sizes. This procedure resulted in $27$ departments and $23$ countries and regions. At the departmental level, we evaluated the proportion of articles originating in benchwork, terming this measure the benchwork score. Lower values corresponded to computational settings, whereas larger values corresponded to research performed predominantly in wet labs. We measured this score using a base version (un-finetuned) of GPT-3.5-Turbo-0125. For each of the $28$ departments, we first sampled $100$ articles, excluding reviews. We then prompted the model to determine whether a given article was a product of benchwork by examining most of its content. The content was extracted using the same PMC parser [154] as the one for title extraction, and up to $10000$ tokens worth of text were included for each article. The benchwork score for a department was defined as the fraction of papers identified by the model having originated in a wet lab. The prompt text supplied to the LLM is available in the Supplement. We also measured synthesis as the proportion of review papers among all papers in a discipline. We identified a review paper by the article-type variable from the XML file for each paper. At the broader cultural scale, we used the 6-D model of national culture to measure power-distance and individualism. Power-distance is the “extent to which the less powerful members of organizations and institutions (like the family) accept and expect that power is distributed unequally.” Individualism, in turn, is the “extent to which people feel independent, as opposed to being interdependent as members of larger wholes.” Values for these measures are available at [79, 66].
VI Citation Diversity Statement
Recent work in several fields of science has identified a bias in citation practices such that papers from women and other minority scholars are under-cited relative to the number of such papers in the field [155, 156, 89, 157, 158, 159, 160, 86, 161]. Here we sought to proactively consider choosing references that reflect the diversity of the field in thought, form of contribution, gender, race, ethnicity, and other factors. First, we obtained the predicted gender of the first and last author of each reference by using databases that store the probability of a first name being carried by a woman [158, 162]. By this measure (and excluding self-citations to the first and last authors of our current paper), our references contain 23.87% woman(first)/woman(last), 12.93% man/woman, 11.81% woman/man, and 51.39% man/man. This method is limited in that a) names, pronouns, and social media profiles used to construct the databases may not, in every case, be indicative of gender identity and b) it cannot account for intersex, non-binary, or transgender people. Second, we obtained predicted racial/ethnic category of the first and last author of each reference by databases that store the probability of a first and last name being carried by an author of color [163, 164]. By this measure (and excluding self-citations), our references contain 14.29% author of color (first)/author of color(last), 16.33% white author/author of color, 15.10% author of color/white author, and 54.28% white author/white author. This method is limited in that a) names and Florida Voter Data to make the predictions may not be indicative of racial/ethnic identity, and b) it cannot account for Indigenous and mixed-race authors, or those who may face differential biases due to the ambiguous racialization or ethnicization of their names. We look forward to future work that could help us to better understand how to support equitable practices in science.
@@ Acknowledgements.
The authors gratefully acknowledge helpful discussions with Drs. Jordan D. Dworkin, Julia K. Brynildsen, and David M. Lydon-Staley.